[Hallc_running] [New Logentry] Hall C RC update Wednesday Sept 5
mack at jlab.org
mack at jlab.org
Wed Sep 5 10:00:03 EDT 2018
As per Arne's guidance, we are cancelling shifts thru Wed Sept 12, 08:00 .
Leak checking in the NL continues. As for Hall C ongoing:
i. The target group declared victory on the fan stall issue after the multi-phase voltage was dropped last Friday. https://logbooks.jlab.org/entry/3593202 This is consistent with fan trip observations during Q-weak, so it's time to move on. Before I leave this topic, I note that the manual for the AC motor controller mentions a +-10% input voltage range, so it's not picky about what it eats. But the relatively high, 3-phase voltage ("480"-ish V) voltage it uses is not something we regularly monitor yet.
ii. Steve Lassiter still plans to move theta_min blocks for the HMS. Given that we're down for (at least) another week, there may be time to hook up the brakes on the motors if Joe Beaufait can arrange electrician support.
iii. No update on whether the wrong software angles would completely explain the vertical motion seen in the big BPM. However, Jay Benesch is confident the operators will be able to find the dump viewer, which is Plan B for the SHMS steering test.
Simona will take over as RC today at 1:30pm . Note there is no Wed scheduling meeting today. There is also no RC meeting today. But if there were, the information to join the meeting would be here:
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3593563
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running